OpenStudy (anonymous):

how to find the maximum y value of -16x^2+32x+100

OpenStudy (anonymous):

max is the second coordinate of the vertex first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \(-\frac{32}{2\times (-16)}=1\) second coordinate is what you get if you replace \(x\) by \(1\)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw